Neglecting air resistance, how long shall it take an object in free fall to have a velocity of 5.0 m/s?


1
Expert's answer
2021-11-16T11:25:07-0500

Explanations & Calculations


  • Since gravitational acceleration is almost constant in low heights where most of the things happen you can use the 4 equations of motion appropriately.
  • The question is asking for the travelled distance until it attains the given velocity having given only the gravitational acceleration and the final velocity you can use "\small v^2=u^2+2as".
  • Then,

"\qquad\qquad
\begin{aligned}
\small h&=\small \frac{v^2-u^2}{2g}\\
&=\small \frac{(5ms^{-2})^2-u^2}{2\times9.8ms^{-2}}

\end{aligned}"

  • If the object falls from rest (u=0) it falls through a distance "\small h=1.3\,m"
  • If it has an initial velocity (u>0) then its velocity needs to be calculated accordingly using that equation.